A nation has a working-age population of 200 million, with 14 million not working and not looking for work and 170 million employed. Find the unemployment rate by writing out the formula and working it out
Answer #1

Employed = 170 million
Unemployed = Total population - Employed - Not looking for work = 200 - 170 - 14 = 16 million

So, labor force = Employed + Unemployed = 170 + 16 = 186 million

Unemployment rate = (Unemployed/Labor force)*100 = (16/186)*100 = 8.6%

Thus, unemployment rate = 8.6%